OpenStudy (chyanne):

Find the quotient. (6x 2 + 23x + 20) ÷ (5 + 2x)

OpenStudy (anonymous):

Notice \[ 6 x^2+23 x+20=(2 x+5) (3 x+4) \]

OpenStudy (anonymous):

Hence \[ \frac {6 x^2+23 x+20}{ 5 + 2 x}=\frac {(2 x+5) (3 x+4)}{ 5 + 2 x}= 3 x +4 \] If x not equal -5/2